OpenStudy (anonymous):

a certain number decreased by 7 multiplied by the sum of the number and 7.the product is 51.find the number.

OpenStudy (anonymous):

(x-7)(x+7)=51 x^2-49=51 x^2=100 x=10 V x=-10

OpenStudy (saifoo.khan):

Yeah he is correct.

OpenStudy (anonymous):

explain..how did you answer it all!!

OpenStudy (saifoo.khan):

no, u r correct.

OpenStudy (anonymous):

Let's call the number we want to find x. Then from the text we find the equation: (x-7)(x+7)=51 writing out the product gives: x^2-49=51 now let's get the 49 on the other side: x^2=100 so x=10 or x=-10. Always be wary of the fact that quadratic equations have two solutions.

OpenStudy (anonymous):

can you check it?

OpenStudy (saifoo.khan):

HE is correct.

OpenStudy (saifoo.khan):

I cant C a good answer here. He owes a Medal.

OpenStudy (anonymous):

check if its correct..

OpenStudy (saifoo.khan):

Medals for every1. ^.^

OpenStudy (anonymous):

10-7=3, 10+7=17 3*17=51 for -10: -10-7=-17, -10+7=-3 -17*-3=51

OpenStudy (anonymous):

ok..thank you so much!!